OpenStudy (anonymous):

If a ≠ b and 1/x + 1/a = 1/b , then x = do any one know how to solve this it would be great

OpenStudy (anonymous):

\[\begin{align*} \frac{1}{x}+\frac{1}{a}&=\frac{1}{b}\\\\ \frac{1}{x}&=\frac{1}{b}-\frac{1}{a}\\\\ \frac{1}{x}&=\frac{a}{ab}-\frac{b}{ab}\\\\ \frac{1}{x}&=\frac{a-b}{ab}\\\\ x&=\frac{ab}{a-b} \end{align*}\]